political polling organization wishes to select a smaller focus group from a group of of 7 Republicans, 3 Democrats, and 4 Independents. In how many ways can the group be chosen if:
a) it will consist of 4 Republicans and 3 Democrats?
b) it will consist of 2 Republicans, 2 Democrats, and 3 Independents?

Answer by stanbon(75887)

political polling organization wishes to select a smaller focus group from a group of of 7 Republicans, 3 Democrats, and 4 Independents. In how many ways can the group be chosen if:
a) it will consist of 4 Republicans and 3 Democrats?
Ans:: 7C4 * 3C3 = (7*6*5)/(1*2*3) *1 = 35 ways
-----------------------------
b) it will consist of 2 Republicans, 2 Democrats, and 3 Independents?
Ans:: 7C2 * 3C2 * 4C3 = 21 * 3 * 4 = 252 ways
=============
Cheers,
Stan H.
==============